# Classical Combinatorial Optimization Scaling for Random Ising Models on 2D Heavy-Hex Graphs

### v1

@@ 3 Results
Figure 2 is an initial and simple way to probe the energy spectrum of these problem instances - in particular we see that there is concentration of the random sampling energy near an approximation ratio of 00 as the system size increases. This means that the problem instances are not incredibly easy to solve in the sense that randomly sampling does not give a large approximation ratio as the system size grows.
Figure 3 shows the Gurobi runtime scaling as a function of system size for the problem instances with no higher-order terms. Figure 4 shows the Gurobi runtime scaling as a function of system size for the problem instances with cubic terms. Figure 4 shows clearly that the relatively un-optimized dimod order-reduction method which introduces many auxiliary terms results in significantly more total computation time. Both Figure 3 and Figure 4 show approximately linear time scaling. The average case runtime in Figures 3 and 4 does not substantially deviate from the largest or smallest runtimes, suggesting that these measured runtimes are reasonably consistent across different random polynomial coefficient choices.
@@
Table 1 quantifies least squares fitting coefficients for four possible types of scaling functions namely logarithmic, linear, quadratic, and exponential. How well these fitted functions model the measured computation time data is quantified in terms of root-mean-square-error (RMSE). The lowest RMSE for all Gurobi runtime scaling are linear and quadratic function fits, for the problem instances solved using both problem encoding methods of ILP and IQP. The lowest RMSE for the Time to Solution at $99\%$ confidence (TTS0.99) using simulated annealing sampling is the exponential function. The TTS exponential curve fits are weakly exponential, with exponents of $0.011$ and $0.0038$ for $100$ and $1000$ variable update sweeps, respectively.
@@
Having the capability to obtain verified optimal ground-states of these optimization problems, using Gurobi, means that we can examine some properties of these ground-states, such as net magnetization. Figure 6 shows the average $Z$-basis magnetization of the optimal solutions found by Gurobi, which shows a slight negative net magnetization, which becomes more pronounced as the system size increases, and is consistent for the problem instances with and without geometrically local cubic terms. However, these plots do not show how consistent this property is for different degenerate ground-states of these systems, or if this is a property of the optimization techniques used by Gurobi. This net magnetization of the ground-state could be a property of this class of Ising models, due to their structure and coefficient distribution, or this could be a result of the algorithms used by the Gurobi software.
4 Discussion and Conclusion
@@
Because this class of Ising models are incredibly sparse they are unsurprisingly relatively computationally easy for classical algorithms. The fundamental motivation for these optimization problems is NISQ hardware-compatibility for a specific type of Quantum Processing Unit (QPU), which allows for scalable near-term quantum computer and algorithmic benchmarking of QAOA [14, 15, 16]. However, in order to find optimization problems that are more computationally challenging for classical algorithms, hardware-compatible problem instances such as those examined in this study will need to include more complex interactions – such as a longer range variable interactions and larger higher order terms that are not necessarily geometrically local. The findings of the results in this study contextualizes where quantum algorithms for combinatorial optimization problems need to get to in order to be competitive with state of the art classical algorithms.
@@ Acknowledgments
Thank you to Carleton Coffrin for discussions on integer programming and Gurobi. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented in this article was supported by the NNSA’s Advanced Simulation and Computing Beyond Moore’s Law Program at Los Alamos National Laboratory. This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). LANL report number LA-UR-24-32607.

### v2

@@ III Results
Figure 2 is an initial and simple way to probe the energy spectrum of these problem instances - in particular we see that there is concentration of the random sampling energy near an approximation ratio of $0.5$ as the system size increases. This means that the problem instances are not trivially easy to solve in the sense that randomly sampling does not give a large approximation ratio as the system size grows.
Figure 3 shows the Gurobi runtime scaling as a function of system size for the problem instances with no higher-order terms, where we see not only very low absolute runtimes but also a scaling that appears to be approximately linear. For these Ising models with no cubic terms, since the binary IQP approach already works very well at finding the optimal solution, we do not additionally run the binary ILP method with Gurobi.
Figure 4 shows the Gurobi runtime scaling as a function of system size for the problem instances with cubic terms. Figure 4 shows clearly that the order-reduction binary IQP method requires more computation time compared to the binary ILP method – a potential reason for this could be the addition of auxiliary variables in the binary IQP case. Both Figure 3 and Figure 4 show approximately linear time scaling. The average case runtime in Figures 3 and 4 does not substantially deviate from the largest or smallest runtimes, suggesting that these measured runtimes are reasonably consistent across different random polynomial coefficient choices.
@@
Table 1 quantifies least squares fitting coefficients for four possible types of scaling functions namely logarithmic, linear, quadratic, and exponential. How well these fitted functions model the measured computation time data is quantified in terms of root-mean-square-error (RMSE). The lowest RMSE for all Gurobi runtime scaling are linear and quadratic function fits, for the problem instances solved using both problem encoding methods of ILP and IQP. The lowest RMSE for the Time to Solution at $99\%$ confidence (TTS0.99) using simulated annealing sampling is the exponential function. The simulated annealing TTS exponential curve fits are weakly exponential, with exponents of $0.00898$ and $0.002$, for $100$ and $1000$ variable update sweeps respectively.
Figure 5: Simulated annealing TTS scaling (log scale y-axis) as a function of Ising model problem size (x-axis) for the instances containing no higher order terms. The x-axis (problem size) is exactly the number of nodes in the heavy-hex graph. $100$ Metropolis-Hastings updates (left) and $1000$ Metropolis-Hastings updates (right). At each system size N, $10$ different random-coefficient instances are solved and then the average time to solve those $10$ instances are shown as a function of $N$ by the connected green line. Here, $N$ refers specifically to the number of nodes in the underlying heavy-hex graph (in other words, not including any auxiliary variables produced from the order reduction process). TTS is measured using $50,000$ (independently generated) samples produced for each problem instance.
@@
Having the capability to obtain verified optimal ground-states of these optimization problems, using Gurobi, means that we can examine some properties of these ground-states, such as net magnetization. Magnetization is a very simple magnetic order parameter, which serves as a simple observable test case that shows that we can use a commercial optimization solver to measure ground-state properties of this class of hardware-defined spin glass models. More complex ground-state order parameters could be extracted from these types of hardware-compatible Ising models, but we leave this to future work. Figure 6 shows the average $Z$-basis magnetization of the optimal solution found by Gurobi [^3] Note that here we are measuring the net magnetization of the single optimal solution found, per problem instance, found by the Gurobi software., which shows a slight negative net magnetization, which becomes more concentrated as the system size increases, and is consistent for the problem instances with and without geometrically local cubic terms. Net magnetization of approximately zero is consistent with this class of Ising models, that being in general ±$\pm J$ models since we have approximately equal contributions from both antiferromagnetic and ferromagnetic interactions. In particular it means that the ground state does not have a very strongly preferred magnetic direction, but a slight ferromagnetic bias. Note that these plots do not show how consistent this property is for different degenerate ground-states of these systems. This net magnetization of the ground-state could be a property of this class of Ising models, due to their structure and coefficient distribution, or this could be a result of the algorithms used by the Gurobi software.
IV Discussion and Conclusion
@@
Primarily because this class of Ising models are incredibly sparse, and quite structured, they are relatively computationally easy for classical algorithms such as commercial mathematical programming software. There are ways that problems despite being very sparse can be computationally challenging, but at least for these classes of problems that this study examines this is not the case – part of this could be due to the geometric locality of the higher order terms, and also potentially because of the low-coefficient precision of the problem instances. The fundamental motivation for these optimization problems is NISQ hardware-compatibility for a specific type of Quantum Processing Unit (QPU), which allows for scalable near-term quantum computer and algorithmic benchmarking of QAOA [14, 15, 16]. However, in order to find optimization problems that are more computationally challenging for classical algorithms (and therefore a potential case where quantum algorithms could yield substantial improvements), promising optimization problem instances will need to include more complex interactions thus moving away from being purely hardware-compatible. Examples of more complex optimization types include longer range variable interactions and higher-weight high-order terms that are not necessarily geometrically local. The findings of the results in this study contextualizes where quantum algorithms performance for solving (whether approximately or exactly) combinatorial optimization problems need to get to in order to be competitive with state of the art classical algorithms, albeit for a specific type of NISQ processor hardware graph. More complexly connected optimization problems also mean that with current noisy quantum processors the degree of meaningful signal that can be extracted is not expected to be very high (as opposed to very hardware compatible optimization problems) [16, 15, 11, 13, 3]. Therefore, in the near term, the central question is whether there exist interesting optimization problems that are both relatively NISQ hardware compatible, but also challenging for state of the art classical optimization solvers. It seems very likely that, if such problems exist, they will need to be on the scale of hundreds of decision variables since problems smaller than that are typically solved very efficiently by existing algorithms.
@@
Importantly, the variable types that are considered in this study being discrete, and in particular spins, are well suited for digital quantum algorithms such as QAOA. And therefore, in future comparisons of quantum algorithms and classical algorithms, it will be necessary to compare against solvers that handle this variable type. However, it is worth noting that many natural types of optimization problems do not inherently have decision variable types that are binary or spins, and in such cases casting to something like an Ising model can be detrimental for the performance of linear programming solvers. In such cases, fair comparisons should be made against state of the art linear programming solvers by not going through extensive problem re-formulation, or order reductions, when utilizing the mathematical optimization software.
@@ Acknowledgments
We thank Carleton Coffrin for discussions on integer programming and Gurobi. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented in this article was supported by the NNSA’s Advanced Simulation and Computing Beyond Moore’s Law Program at Los Alamos National Laboratory. This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE).
LA-UR-24-32607